# Beyond Numeric Awards: In-Context Dueling Bandits with LLM Agents

### v3

@@ Abstract
In-context reinforcement learning (ICRL) is a frontier paradigm for solving reinforcement learning problems in the foundation model era. While ICRL capabilities have been demonstrated in transformers through task-specific training, the potential of Large Language Models (LLMs) out-of-the-box remains largely unexplored. Recent findings highlight that LLMs often face challenges when dealing with numerical contexts, and limited attention has been paid to evaluating their performance through preference feedback generated by the environment. This paper is the first to investigate LLMs as in-context decision-makers under the problem of Dueling Bandits (DB), a stateless preference-based reinforcement learning setting that extends the classic Multi-Armed Bandit (MAB) model by querying for preference feedback. We compare GPT-3.5 Turbo, GPT-4, GPT-4 Turbo, Llama 3.1, and o1-preview against nine well-established DB algorithms. Our results reveal that our top-performing LLM, GPT-4 Turbo, has the zero-shot relative decision-making ability to achieve surprisingly low short-term weak regret across all the DB environment instances by quickly including the best arm in duels. However, an optimality gap exists between LLMs and classic DB algorithms in terms of strong regret. LLMs struggle to converge and consistently exploit even when explicitly prompted to do so, and are sensitive to prompt variations. To bridge this gap, we propose an agentic flow framework: LLM with Enhanced Algorithmic Dueling (LEAD), which integrates off-the-shelf DB algorithms with LLM agents through fine-grained adaptive interplay. We show that LEAD has theoretical guarantees inherited from classic DB algorithms on both weak and strong regret. We validate its efficacy and robustness even with noisy and adversarial prompts. The design of such an agentic framework sheds light on how to enhance the trustworthiness of general-purpose LLMs used for in-context decision-making tasks.
@@ 1 Introduction
Large sequence models pretrained with task specific interactive datasets have led to the emergence of in-context reinforcement learning (ICRL) (Laskin et al., 2022; Lee et al., 2024), where models can infer tasks from interaction histories as context and make effective decisions in unseen environments without parameter updates. Through trial and error, these models can self-improve their policies purely in-context. While ICRL capabilities have been demonstrated in transformers with task-specific training from scratch, the potential of general-purpose Large Language Models (LLMs) to perform ICRL remains largely unexplored. Recent investigations into LLMs’ out-of-the-box ICRL capabilities in environments with numeric rewards have reported notable failure cases, e.g., LLM agents being vulnerable to adversarial loss functions and suffering from high regret compared to classic algorithms such as Follow-The-Regularized-Leader (FTRL) (Park et al., 2024), and exhibiting failures in exploration within Multi-Armed Bandit (MAB) problems via standard training (Krishnamurthy et al., 2024). Even with inference-time algorithmic guidance, an optimality gap persists between LLMs and classic (contextual) MAB algorithms (Nie et al., 2024). These results suggest that carefully designed prompts and non-trivial algorithmic interventions are needed to elicit desirable in-context reinforcement learning behavior of LLM agents.
The failure cases encountered by LLMs may be attributed to intrinsic difficulties in processing numeric rewards, especially in tasks where patterns are difficult to express in natural language. Recent findings have pointed out that LLMs often struggle with simple numerical comparisons (e.g., incorrectly judging 13.11 to be larger than 13.8), and there has been a notable lack of emphasis on evaluating the relative comparisons among the decisions they generate. Figure 1 shows a toy example illustrating the in-context interaction between an LLM agent and different environment settings. To disentangle the complexities introduced by numerical rewards, this paper focuses on the problem of Dueling Bandits (DB) (Yue et al., 2012; Zoghi et al., 2014a), a stateless preference-based reinforcement learning setting (Wirth et al., 2017; Pacchiano et al., 2021) that extends the classic MAB model by querying for preference feedback between selected pairs of arms to identify the best one. In DB, the agent learns through binary outcome (win or lose) of a noisy comparison between the two selected arms. This setup is particularly useful when eliciting explicit feedback is challenging or when the feedback is inherently comparative, like taste of food and product attractiveness (Yue et al., 2012). DB has attracted significant attention due to its applicability in information retrieval (Yue and Joachims, 2009), recommendation systems (Sui et al., 2017), and online ranker evaluation (Zoghi et al., 2014a). We frame our investigation with the following question:
@@
The DB problem poses distinctive challenges as a relative decision-making instance, particularly due to the sparse nature of the relative rewards. This sparsity complicates the in-context decision-making process, as it restricts the feedback obtained from interactions, introducing a level of difficulty not typically seen in conventional bandit problems. Even though reduction from DB to standard MAB exists (Ailon et al., 2014; Saha and Gaillard, 2022), it remains unclear how LLMs would perform in DB with preference feedback rather than numeric rewards. There are conceptual differences between them, similar to those between Reinforcement Learning from Human Feedback (RLHF) (Stiennon et al., 2020) and standard RL, where impossibility results can be found in (Wang et al., 2024a).
@@
While task-specific training of large sequence models can yield promising ICRL results, it is often impractical due to the substantial computational resources required. Similar to the settings in (Krishnamurthy et al., 2024; Nie et al., 2024; Mirchandani et al., 2023; Chen et al., 2024), we evaluate the emergent zero-shot abilities (Wei et al., 2022) of ICRL in general-purpose LLMs under the dueling bandit problem, without re-training or fine-tuning. We summarize our main results below.
Figure 1: In-context reinforcement learning of an LLM agent with numeric rewards (in a multi-armed bandit environment) and preference feedback (in a dueling bandit environment).
@@
Evaluation of LLMs’ Emergent Zero-Shot Abilities of In-Context Dueling Bandits. We go beyond numeric rewards to evaluate the performance of LLM agents in terms of both strong and weak regret for making decisions in DB by comparing against various baseline DB algorithms via a case study. We found that the top-performing general-purpose LLMs has the zero-shot ability for relative decision-making sufficient to achieve low weak regret in DB, which significantly differs from that in classic MAB settings (Krishnamurthy et al., 2024). Notably, GPT-4 Turbo can serve as an effective decision-maker for dueling bandits in terms of weak regret, quickly selecting the best arm in duels with low variance across a range of instances. However, consistent with (Nie et al., 2024), we found that an optimality gap exists between LLMs and classic DB algorithms in terms of strong regret. LLMs’ performance is hindered by over-estimation bias in the exploration stage and lack of convergence criterion in the exploitation stage. This highlights the need for more effective and robust strategies to bridge this gap for in-context dueling bandits.
@@
Effective and Robust Agentic Flow Framework for In-Context Dueling Bandits. To address the identified optimality gap and enhance the trustworthiness of in-context LLM agents in DB tasks, in Section 4.1, we propose an agentic flow framework, LLM with Enhanced Algorithmic Dueling (LEAD) that integrates off-the-shelf Explore-then-Exploit DB algorithms with LLM agents. This framework enables the fine-grained adaptive interplay between rule-based expert systems and in-context LLM agents, enhancing their ability to handle DB problems via algorithmic interventions as suggested in (Krishnamurthy et al., 2024; Nie et al., 2024). As an illustrative example, we demonstrate how Interleaved Filter2 (IF2) algorithm can be incorporated with LLM agents in this framework. We show that LEAD has theoretical guarantees, with experiments demonstrating its efficacy and robustness across various prompting scenarios.
@@ 2 Preliminaries
Dueling Bandits. In a fundamental $K$-armed dueling bandit problem setting (Yue et al., 2012), a learner interacts with the environment by selecting two arms $\mathsf{Arm}_{1}(t)$ and $\mathsf{Arm}_{2}(t)$ from a set of $K$ arms $\{b_{1},\ldots,b_{K}\}$ for a noisy comparison (a duel), at each round $t\in\{1,\ldots,T\}$ as Figure 1 illustrates. The outcome of a duel between two arms $(i,j)$ is probabilistic. More precisely, the event that an arm $b_{i}$ wins against $b_{j}$ is a Bernoulli random variable with a parameter denoted by $\Pr(b_{i}\succ b_{j})$. For notational convenience, we normalize $\Pr(b_{i}\succ b_{j})$ such that $\Pr(b_{i}\succ b_{j})=\epsilon(b_{i},b_{j})+1/2$, where $\epsilon_{ij}\coloneq\epsilon(b_{i},b_{j})\in(-1/2,1/2)$ is a measure of the distinguishability between arms $b_{i}$ and $b_{j}$, which is stationary over time and is symmetric such that $\epsilon_{ij}=-\epsilon_{ji}$ for all $i,j\in[K]\coloneq\{1,\ldots,K\}$. Finally, for notational convenience, we define a preference matrix $P=\left[\epsilon_{ij}\right]_{i,j\in[K]}$.
In-Context LLM Agents for Dueling Bandits. We consider an LLM agent with policy $\pi_{\mathrm{LLM}}$ interacting with a $K$-armed dueling bandit environment in-context. At each round $t\in\{1,\ldots,T\}$, the LLM agent selects a pair of arms $(\mathsf{Arm}_{1}(t),\mathsf{Arm}_{2}(t))$ from the set $\{b_{1},\ldots,b_{K}\}$ based on a natural language instruction $\mathtt{Prompt}(C,H_{t},R)$ (see Figure 7), consisting of three parts:
@@
  • History $H_{t}$: an externally summarized interaction history (Krishnamurthy et al., 2024) up to round $t$, which includes a sequence of pairwise dueling results and the empirical probabilities.
@@
$\left(\mathsf{Arm}_{1}(t),\mathsf{Arm}_{2}(t)\right)=\pi_{\mathrm{LLM}}\left(\mathtt{Prompt}(P,H_{t},R)\right).$ (1)
@@
Strong and Weak Regret. Throughout this paper, we assume the standard setting that a Condorcet winner (CW) exists (Sui et al., 2017; Wu and Liu, 2016; Zoghi et al., 2014a; Yue et al., 2012). The CW denoted as $b^{*}$ is an arm that is preferred over all the other arms, i.e., $b^{*}=b_{i}$ if $\epsilon_{ij}>1/2$ for all $j\in[K]\backslash\{i\}$. We consider two performance metrics: (i) strong regret (SR), which evaluates the total preference gap between $b^{*}$ and both selected arms; (ii) weak regret (WR), which compares $b^{*}$ only with the better of the two arms. Detailed definitions and settings are provided in Appendix B.3.1.
@@ 3 LLMs as Standalone In-Context Decision-Makers
3.1 Implementation Details of Experiments
Prompts and Configurations of LLMs. We employ an interactive zero-shot chain-of-thought (CoT) prompt $\mathtt{Prompt}(P,H_{t},R)$, as defined in Section 2, which describes the problem setting $P$, externally summarized interaction history $H_{t}$ and reasoning instructions $R$. We adopt the prompting template and LLM configurations that lead to the best performance among all prompt variations explored in recent studies (Krishnamurthy et al., 2024; Nie et al., 2024) for MAB problem. The LLM agents interact with dueling bandit environments in a round-based manner, with the prompt guiding their decision-making process. We conduct experiments with five LLMs: GPT-3.5 Turbo, GPT-4, GPT-4 Turbo, Llama 3.1, and o1-preview. Note that we skip the GPT-4o version which is primarily developed for multimodal tasks and has the same intelligence as GPT-4 Turbo. The detailed prompt is provided in Appendix C.1.2.
Baselines. We compare LLMs against nine well-established baseline algorithms to evaluate their efficacy. The baselines include Interleaved Filter (IF2(Yue et al., 2012), Beat the Mean (BTM(Yue and Joachims, 2011), Sensitivity Analysis of VAriables for Generic Exploration (SAVAGE(Urvoy et al., 2013), Relative Upper Confidence Bound (RUCB(Zoghi et al., 2014b), Relative Confidence Sampling (RCS(Zoghi et al., 2014a), Relative Minimum Empirical Divergence (RMED(Komiyama et al., 2015), Versatile Dueling Bandits (VDB(Saha and Gaillard, 2022), Self-Sparring (Sui et al., 2017), and Double Thompson Sampling (DTS(Wu and Liu, 2016). Each of these algorithms employs distinct strategies for selecting arms and estimating preferences, with the ultimate goal of efficiently identifying the Condorcet winner. We assess the performance of LLMs and baseline algorithms using strong regret and weak regret metrics defined in Section 2.
[…]
@@ 3.2 Experimental results
For brevity, we present our initial analysis focused on the Transitive-Easy instance (Figure 2). The analysis is qualitatively similar for the Transitive-Hard instance (Figure 9 in Appendix C.2). We use $\gamma=0.5$ for BTM, $f(K)=0.3K^{1.01}$ for RMED, $\eta=1$ for Self-Sparring, and $\alpha=0.51$ for RUCB, RCS and DTS. We analyze the results in terms of the strong and weak regret defined in Section 2. In the following sections, we will mainly focus on GPT-4 Turbo, which is our top-performing LLM, highlighting its success and failure modes.
Emergence of in-context dueling bandits abilities. While GPT-3.5 Turbo and GPT-4 fail to solve the DB problem, GPT-4 Turbo consistently outperforms state-of-the-art DB baselines in weak regret on Transitive Case (see Figures 2 and 9). This reveals that the in-context dueling bandits abilities emerge as the general capabilities grow via standard training in general-purpose LLMs. Figure 13 (Left) illustrates the fraction of duels including the best arm across different time intervals. GPT-4 Turbo outperforms other LLMs and the DB baselines throughout the entire timeline. These findings suggest that GPT-4 Turbo can effectively process the preference feedback obtained from duels and make informed decisions to quickly identify and include the best arm in its duels.
@@
Best Arm Identification: LLMs’ in-context dueling bandits abilities emerge as the general capabilities grow. The Condorcet Winner is consistently selected in duel via GPT-4 Turbo, leading to exceptional weak regret performance with minimal variance on Transitive Case.
Exploration Vulnerability. In the exploration stage, we observe that GPT-4 Turbo tends to quickly narrow down to a small subset of arms (although usually containing the Condorcet Winner) and repeatedly compare these arms. In contrast, the baselines exhibit more diverse and extended exploration patterns. This behavior suggests that GPT-4 Turbo may overestimate the quality of arms that win their initial comparisons based on limited historical data. Unlike the baselines, which have explicit exploration mechanisms, LLMs rely on their inherent randomness (via sampling from their output distribution) to explore. Based on these observations, we hypothesize that if GPT-4 Turbo happens to sample a sequence of comparisons that favors suboptimal arms early on, it can get stuck comparing these arms indefinitely. To test this hypothesis, we conducted experiments using noisy prompts with biased history. Our results in Figure 16 confirm that GPT-4 Turbo’s exploration strategy is indeed vulnerable to biased history initialization and can converge to local optima.
Exploitation Inability. Despite GPT-4 Turbo’s outstanding weak regret performance, it fails to consistently converge to a single best arm to duel against itself, even when the prompt setting explicitly calls for it. This behavior highlights a fundamental limitation of LLMs: they are primarily designed and trained for word token prediction rather than decision-making. Unlike baselines with explicit stopping conditions, GPT-4 Turbo relies on its inherent language modeling capabilities to determine when to stop exploring. Consequently, in the later exploitation stage, GPT-4 Turbo keeps comparing the same top arms without committing to a single winner (see Figure 3). This suggests that the language modeling objective alone may not be sufficient for LLMs to achieve optimal control in complex decision-making tasks like DB.
@@
Biased understanding of DB problem during pretraining. Our two best-performing LLMs, GPT-4 Turbo and o1-preview, exhibit systematic biases regarding the DB problem, likely due to a lack of exposure to similar tasks during pretraining. Specifically, they incorrectly assume that an arm cannot duel with itself (the convergence case), even when explicitly prompted to do so (see examples in Appendix C.1.3). This misunderstanding makes the DB problem as an out-of-distribution (OOD) task for LLMs, and in-context instructions fail to fully override this internal bias. Consequently, LLM agents cannot completely align with problem descriptions due to the inherent limitations of in-context learning, which cannot really generalize to OOD tasks (Wang et al., 2024b). Figure 13 supports these observations: o1-preview demonstrates better reasoning capabilities by transitioning from exploration to exploitation effectively and achieving lower strong regret than GPT-4 Turbo. However, its inference-time CoT mechanism reinforces its internal biased understanding of DB, resulting in bad weak regret performance due to the selection of two suboptimal arms in duels.
@@
Scalability Limitation. To evaluate whether LLMs can generalize their exceptional weak regret performance, we conduct experiments from two perspectives: (i) Removing transitivity in preference structures: we change from transitive cases to intransitive cases that violate SST and STI (see Figures 10 and 11). The analysis of Intransitive-Easy and Intransitive-Hard is qualitatively similar: LLMs fail to replicate their long-term weak regret performance in transitive cases when faced with intransitive instances. However, their short-term weak regret performance (< 100 steps) remains exceptional; (ii) Increasing the number of arms: as illustrated in Figure 12, from $K=5$ to $K=10$, GPT-4 Turbo’s performance exhibits a noticeable long-term decline with the increase in $K$. While LLMs still beat all the DB baselines in the initial steps, they struggle to effectively infer the relative strengths among a larger number of arms in the long run. These findings suggest that while LLMs exhibit emergent abilities for relative decision-making rooted in linguistic knowledge, their effectiveness is only generalized to short-term scenarios. The long-term performance is hindered by larger number of arms and LLMs’ lack of fundamental understanding of intransitive cyclic structures.
To quantify this long-term scalability limitation and formally characterize the dueling bandit instances LLMs can handle, we introduce the concept of the Relative Decision Boundary (RDB). The RDB for a given LLM $m$ is defined as the set of problem difficulties $D$ for which the model achieves an acceptable level of weak regret across an effective short-term time horizon $T_{e}$, satisfying the condition:
$\text{RDB}(m)=\left\{(K,T,\Delta,T_{e})\ \Big{|}\ \mathsf{WR}(m,D(K,T,\Delta),T_{e})\leq\mathsf{R_{\text{th}}}\right\}.$ (2)
Here, $\mathsf{WR}(m,D,T_{e})$ represents the cumulative weak regret incurred by the model $m$ on a problem of difficulty $D$ over an effective short-term time horizon $T_{e}$, while $\mathsf{R_{\text{th}}}$ is a predefined threshold that quantifies acceptable weak regret performance. Overall, RDB is influenced by the inherent ability of $m$, the number of arms $K$, the transitivity $T$, the distinguishability between arms $\Delta$, and the effective short-term time horizon $T_{e}$.
Failure of long-term generalization: LLMs’ long-term strong and weak regret performance degrades when introducing intransitive preference structures or large number of arms. Success of short-term generalization: LLMs’ short-term weak regret performance remains surprisingly exceptional across all instances (see subfigures in 2,  9101112). We introduce Relative Decision Boundary (RDB) to describe the dueling bandit instances that LLMs can effectively handle.
As a summary, we found that in-context LLM agents’ linguistic prior allows them to quickly identify the Condorcet Winner from the dueling history in the short-term (for both Transitive Case and Intransitive Case), but it is vulnerable.
To further investigate the algorithmic behavior of LLMs and develop more robust and effective in-context decision-making strategies, we seek to answer the following questions:

### v4

# Beyond Numeric Rewards: In-Context Dueling Bandits with LLM Agents

@@ Abstract
In-Context Reinforcement Learning (ICRL) is a frontier paradigm to solve Reinforcement Learning (RL) problems in the foundation model era. While ICRL capabilities have been demonstrated in transformers through task-specific training, the potential of Large Language Models (LLMs) out-of-the-box remains largely unexplored. This paper investigates whether LLMs can generalize cross-domain to perform ICRL under the problem of Dueling Bandits (DB), a stateless preference-based RL setting. We find that the top-performing LLMs exhibit a notable zero-shot capacity for relative decision-making, which translates to low short-term weak regret across all DB environment instances by quickly including the best arm in duels. However, an optimality gap still exists between LLMs and classic DB algorithms in terms of strong regret. LLMs struggle to converge and consistently exploit even when explicitly prompted to do so, and are sensitive to prompt variations. To bridge this gap, we propose an agentic flow framework: LLM with Enhanced Algorithmic Dueling (LEAD), which integrates off-the-shelf DB algorithm support with LLM agents through fine-grained adaptive interplay. We show that LEAD has theoretical guarantees inherited from classic DB algorithms on both weak and strong regret. We validate its efficacy and robustness even with noisy and adversarial prompts. The design of such an agentic framework sheds light on how to enhance the trustworthiness of general-purpose LLMs generalized to in-context decision-making tasks.
[…]
tcb@breakable
Beyond Numeric Rewards: In-Context Dueling Bandits with LLM Agents
@@ 1 Introduction
Transformers pretrained with interactive datasets have led to the emergence of In-Context Reinforcement Learning (ICRL) (Laskin et al., ; Lee et al., 2024), where models can infer Reinforcement Learning (RL) tasks from interaction histories as context and make effective decisions without parameter updates. Through trial and error, these models can self-improve their policies purely in-context. Recent investigations into LLMs’ ICRL capabilities in environments with numeric rewards have reported notable failure cases, e.g., LLM agents being vulnerable to adversarial loss functions and suffering from high regret compared to classic algorithms such as Follow-The-Regularized-Leader (FTRL) (Park et al., ), and exhibiting failures in exploration within Multi-Armed Bandit (MAB) problems (Krishnamurthy et al., ). Even with inference-time algorithm guidance, an optimality gap persists between LLMs and classic MAB algorithms (Nie et al., 2024). These results highlight the need for further research in non-trivial algorithmic interventions to elicit desirable ICRL behavior in LLM agents.
Figure 1: In-context reinforcement learning of an LLM agent in a Multi-Armed Bandit (MAB) environment and a Dueling Bandit (DB) environment.
The failure cases encountered by LLMs may be attributed to intrinsic difficulties in processing numeric rewards, especially in tasks where patterns are difficult to express in natural language. Recent findings have pointed out that LLMs often struggle with simple numerical comparisons (e.g., incorrectly judging 13.11 to be larger than 13.8), and there has been a notable lack of emphasis on evaluating the relative comparisons among the decisions they generate. Figure 1 shows a toy example illustrating the in-context interaction between an LLM agent and different environment settings. To disentangle the complexities introduced by numerical rewards, this paper focuses on the problem of Dueling Bandits (DB) (Yue et al., 2012; Zoghi et al., 2014b), a stateless preference-based reinforcement learning setting (Wirth et al., 2017; Saha et al., 2023) that extends the classic MAB model by querying for preference feedback between selected pairs of arms to identify the best one. In DB, the agent learns through binary outcome (win or lose) of a noisy comparison between the two selected arms. This setup is particularly useful when eliciting explicit feedback is challenging or when the feedback is inherently comparative, such as the taste of food and product attractiveness (Yue et al., 2012). DB has attracted significant attention due to its applicability in information retrieval (Yue and Joachims, 2009), recommendation systems (Sui et al., 2017), and online ranker evaluation (Zoghi et al., 2014b). We frame our investigation with the following question:
@@
The DB problem poses distinctive challenges as a relative decision-making instance, particularly due to the sparse nature of the relative rewards. This sparsity complicates the in-context decision-making process, as it restricts the feedback obtained from interactions, introducing a level of difficulty not typically seen in conventional bandit problems. Even though reduction from DB to standard MAB exists (Ailon et al., 2014; Saha and Gaillard, 2022), it remains unclear how LLMs would perform in DB with preference feedback rather than numeric rewards. There are conceptual differences between them, similar to those between Reinforcement Learning from Human Feedback (RLHF) (Stiennon et al., 2020) and standard RL, where impossibility results can be found in (Wang et al., 2024b).
@@
While task-specific training of large sequence models can yield promising ICRL results, it is often impractical due to the substantial computational resources required. Similar to the settings in (Krishnamurthy et al., ; Nie et al., 2024; Mirchandani et al., 2023; Chen et al., 2024), we evaluate the emergent zero-shot abilities (Wei et al., ) of ICRL in general-purpose LLMs under the dueling bandit problem, without re-training or fine-tuning. We summarize our main results below.
@@
Evaluation of LLMs’ emergent zero-shot abilities of in-context DB. We go beyond numeric rewards to evaluate the performance of LLM agents in terms of both strong and weak regret for making decisions in DB by comparing against various baseline DB algorithms via a case study. We found that our top-performing LLM has the zero-shot relative decision-making ability sufficient to achieve low weak regret in DB, which significantly differs from that in classic MAB settings (Krishnamurthy et al., ). Notably, GPT-4 Turbo can serve as an effective decision-maker for dueling bandits in terms of weak regret, quickly selecting the best arm in duels with low variance across a range of instances. However, consistent with (Nie et al., 2024), we found that an optimality gap exists between LLMs and classic DB algorithms in terms of strong regret. LLMs’ performance is hindered by over-estimation bias in the exploration stage and lack of convergence criterion in the exploitation stage. This highlights the need for non-trivial algorithmic interventions to achieve cross-domain generalization.
@@
Effective and robust agentic flow framework for in-context DB. To address the identified optimality gap and enhance the trustworthiness of in-context LLM agents in DB tasks, in Section 4.1, we propose an agentic flow framework, LLM with Enhanced Algorithmic Dueling (LEAD) that integrates off-the-shelf Explore-then-Exploit DB algorithms with LLM agents. This framework enables the fine-grained adaptive interplay between DB algorithms and in-context LLM agents, pushing forward from the naive algorithm-guided support used in (Nie et al., 2024). As an illustrative example, we demonstrate how the Interleaved Filter2 (IF2) algorithm can be incorporated with LLM agents in this framework. We show that LEAD has theoretical guarantees, with experiments demonstrating its efficacy and robustness across various prompting scenarios.
@@ 2 Preliminaries
Dueling bandits. In a fundamental context-free $K$-armed dueling bandit problem setting (Yue et al., 2012), a learner interacts with the environment by selecting two arms $\mathsf{Arm}_{1}(t)$ and $\mathsf{Arm}_{2}(t)$ from a set of $K$ arms $\{b_{1},\ldots,b_{K}\}$ for a noisy comparison (a duel), at each round $t\in\{1,\ldots,T\}$ as Figure 1 illustrates. The outcome of a duel between two arms $(i,j)$ is probabilistic. More precisely, the event that an arm $b_{i}$ wins against $b_{j}$ is a Bernoulli random variable with a parameter denoted by $\Pr(b_{i}\succ b_{j})$. For notational convenience, we normalize $\Pr(b_{i}\succ b_{j})$ such that $\Pr(b_{i}\succ b_{j})=\epsilon(b_{i},b_{j})+1/2$, where $\epsilon_{ij}\coloneq\epsilon(b_{i},b_{j})\in(-1/2,1/2)$ is a measure of the distinguishability between arms $b_{i}$ and $b_{j}$, which is stationary over time and is symmetric such that $\epsilon_{ij}=-\epsilon_{ji}$ for all $i,j\in[K]\coloneq\{1,\ldots,K\}$. Finally, for notational convenience, we define a preference matrix $P=\left[\epsilon_{ij}\right]_{i,j\in[K]}$.
In-context LLM agents for dueling bandits. We consider an LLM agent with policy $\pi$ interacting with a $K$-armed dueling bandit environment in-context. At each round $t\in\{1,\ldots,T\}$, the LLM agent selects a pair of arms $(\mathsf{Arm}_{1}(t),\mathsf{Arm}_{2}(t))$ from the set $\{b_{1},\ldots,b_{K}\}$ based on a natural language instruction $\mathtt{Prompt}(P,H_{t},R)$ (see Figure 7), consisting of three parts:
@@
  • History $H_{t}$: an externally summarized interaction history (Krishnamurthy et al., ) up to round $t$, which includes a sequence of pairwise dueling results and the empirical probabilities.
@@
$\left(\mathsf{Arm}_{1}(t),\mathsf{Arm}_{2}(t)\right)=\pi\left(\mathtt{Prompt}(P,H_{t},R)\right).$ (1)
@@
Strong and weak regret. Throughout this paper, we assume the standard setting that a Condorcet winner (CW) exists (Sui et al., 2017; Wu and Liu, 2016; Zoghi et al., 2014b; Yue et al., 2012). The CW denoted as $b^{*}$ is an arm that is preferred over all the other arms, i.e., $b^{*}=b_{i}$ if $\epsilon_{ij}>1/2$ for all $j\in[K]\backslash\{i\}$. We consider two performance metrics: (i) strong regret (SR), which evaluates the total preference gap between $b^{*}$ and both selected arms; (ii) weak regret (WR), which compares $b^{*}$ only with the better of the two arms. Detailed definitions and settings are provided in Appendix B.3.1.
@@ 3.1 Experimental results
Implementation details of experiments. In Appendix C, we provide the following implementation details: (i) The LLM configurations and prompting templates; (ii) The baseline algorithms used for comparison; (iii) The DB environments, including transitive and intransitive cases; (iv) The scale setup of our experiments.
For brevity, we present our initial analysis focused on the Transitive-Easy instance (Figure 2). The analysis is qualitatively similar for the Transitive-Hard instance (see Figure 9 in Appendix). We analyze the results in terms of the strong and weak regret defined in Section 2. In the following sections, we will mainly focus on GPT-4 Turbo, which is our top-performing LLM, highlighting its success and failure modes.
Emergence of in-context DB abilities. While GPT-3.5 Turbo and GPT-4 fail to solve the DB problem, GPT-4 Turbo consistently outperforms state-of-the-art DB baselines in weak regret on the Transitive Case (see Figures 2 and 9). This reveals that the in-context DB abilities emerge as the general capabilities grow in general-purpose LLMs. Figure 13 (Left) illustrates the fraction of duels including the best arm across different time intervals. GPT-4 Turbo outperforms other LLMs and the DB baselines throughout the entire timeline. These findings suggest that GPT-4 Turbo can effectively process the preference feedback obtained from duels and make informed decisions to quickly identify and include the best arm in its duels.
@@
Best Arm Identification: LLMs’ in-context dueling bandit abilities emerge as the general capabilities grow. The Condorcet Winner is consistently selected in duels via GPT-4 Turbo, leading to exceptional weak regret performance with minimal variance on the Transitive Case.
Exploration vulnerability. In the exploration stage, we observe that GPT-4 Turbo tends to quickly narrow down to a small subset of arms (although usually containing the Condorcet Winner) and repeatedly compare these arms. In contrast, the baselines exhibit more diverse and explicit exploration patterns. This behavior suggests that GPT-4 Turbo may overestimate the quality of arms that win their initial comparisons based on limited historical data. Based on these observations, we hypothesize that if GPT-4 Turbo happens to sample a sequence of comparisons that favors suboptimal arms early on, it can get stuck comparing these arms indefinitely. To test this hypothesis, we conducted experiments using noisy prompts with biased history. Our results in Figure 16 confirm that GPT-4 Turbo’s exploration strategy is indeed vulnerable to biased history initialization and can converge to local optima.
Exploitation inability. Despite GPT-4 Turbo’s outstanding weak regret performance, it fails to consistently converge to a single best arm to duel against itself, even when the prompt setting explicitly calls for it. Unlike baselines with explicit stopping conditions, GPT-4 Turbo relies on its inherent language modeling capabilities to determine when to stop exploring. Consequently, in the later exploitation stage, GPT-4 Turbo keeps comparing the same top arms without committing to a single winner (see Figure 3). This suggests that the language modeling objective alone may not be sufficient for LLMs to generalize effectively in complex decision-making tasks like DB.
@@
Biased understanding of DB problem during pretraining. Our two best-performing LLMs, GPT-4 Turbo and o1-preview, exhibit systematic biases regarding the DB problem, likely due to a lack of exposure to similar tasks during pretraining. Specifically, they incorrectly assume that an arm cannot duel with itself (the convergence case), even when explicitly prompted to do so (see examples in Appendix C.2.3). This misunderstanding makes the DB problem an out-of-distribution (OOD) task for LLMs, and in-context instructions fail to fully override this internal bias. Consequently, LLM agents cannot completely align with problem descriptions due to the inherent limitations of in-context learning, which cannot really generalize to OOD tasks (Wang et al., 2024a). Figure 13 supports these observations: o1-preview demonstrates better reasoning capabilities by transitioning from exploration to exploitation effectively and achieving lower strong regret than GPT-4 Turbo. However, its inference-time CoT mechanism reinforces its internal biased understanding of DB, resulting in bad weak regret performance due to the selection of two suboptimal arms in duels.
@@
Scalability. To evaluate whether LLMs can generalize their exceptional weak regret performance, we conduct experiments from two perspectives: (i) Removing transitivity in preference structures: we change from transitive cases to intransitive cases that violate SST and STI (see Figures 10 and 11). The analysis of Intransitive-Easy and Intransitive-Hard is qualitatively similar: LLMs fail to replicate their long-term weak regret performance in transitive cases when faced with intransitive instances. However, their short-term weak regret performance remains exceptional; (ii) Increasing the number of arms: as illustrated in Figure 12, from $K=5$ to $K=10$, GPT-4 Turbo’s performance exhibits a noticeable long-term decline with the increase in $K$. While LLMs still beat all the DB baselines in the initial steps, they struggle to effectively infer the relative strengths among a larger number of arms in the long run. These findings suggest that while LLMs exhibit emergent abilities for relative decision-making rooted in linguistic knowledge, their effectiveness only generalizes to short-term scenarios. The long-term performance is hindered by a larger number of arms and LLMs’ lack of fundamental understanding of intransitive cyclic structures.
Success of short-term generalization: Although LLMs’ long-term strong and weak regret performance degrades when introducing intransitive preference structures or a large number of arms, their short-term weak regret performance remains surprisingly exceptional across all instances (see subfigures in 2,  9101112).
As a summary, in-context LLM agents’ linguistic prior allows them to quickly identify the Condorcet Winner from the dueling history in the short term (for both Transitive Case and Intransitive Case), but it is vulnerable. To further investigate the algorithmic behavior of LLMs and develop more robust and effective in-context decision-making strategies, we seek to answer the following questions: